Grade Upto college level Engineering Entrance Exams

A standing wave having 3 nodes and 2 antinodes is found between two atoms having a distance of 1.2A°. between them the wavelength of standing wave will be (a) 3.63 A° (b) 6.05 A° (c) 1.21 A° (d) 2.42 A°

sumit kumar
12 Years agoGrade Upto college level

2 Answers

Arun
6 Years ago
 
Distance between a node and adjoining = λ/4 From figure, distance between two atoms
 
thanks and regards
Arun
Vikas TU
6 Years ago
In the standing wave, if we have 3 nodes, then
Distance between two atoms is given in term of wavelength = λ/2+λ/2 = λ
=1.21A0

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago